# Parametrically enhancing sensor sensitivity at an exceptional point

### 12 Dec 2023 v2

@@ Appendix A Effective mechanical Hamiltonian
$H_{\rm{drive}}$ $=iE\sum_{j}(a^{\dagger}_{j}-a_{j}),$ (29c)
$H_{\rm{diss}}^{S}$ $=-i\frac{\kappa}{2}\sum_{j}a^{\dagger}_{j}a_{j}-i\frac{\gamma_{m}}{2}\sum_{j}b^{\dagger}_{s_{j}}b_{s_{j}},$ (29d)
where the Hamiltonian $H_{\rm{diss}}$ captures the optical ($\kappa$) and the mechanical ($\gamma_{1}=\gamma_{2}=\gamma_{m}$) dissipation in the system. Furthermore, the driving amplitude has been expressed as $E=\sqrt{\kappa}a^{in}$, where $\alpha^{in}$ is related to the input power $P_{in}$ through $\alpha^{in}=\sqrt{\frac{P_{in}}{\hbar\omega_{p}}}$. The following parameters have been also defined,
$\tilde{g}$ $=ge^{-r},$ (30a)
@@
The Quantum Langeving Equations (QLEs) of the Hamiltonian given in Appendix A can be derived using the Heisenberg equation $\dot{\mathcal{O}}=i[H,\mathcal{O}]=\mathcal{N}$, where $H$ is a Hamiltonian, $\mathcal{O}$ is an operator ($\mathcal{O}\equiv b_{s_{j}},a_{j}$), and $\mathcal{N}$ the quantum noise related to the operator. Following this procedure, the QLEs of our system yield,
$\dot{a}_{j}$ $=$ $(i[\Delta_{j}+\tilde{g}(b_{s_{j}}+b^{\dagger}_{s_{j}})]-\frac{\kappa}{2})a_{j}+\sqrt{\kappa}\alpha^{in}_{j}+\sqrt{\kappa}a^{in}_{j},$ (31a)
$\dot{b}_{s_{j}}$ $=$ $-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})b_{s_{j}}+i\tilde{J}_{m}b_{3-j}+i\tilde{g}a^{\dagger}_{j}a_{j}+\sqrt{\gamma_{m}}\beta^{in}_{j}.$ (31b)
The nonlinear set of equations displayed in Appendix A can be linearized by splitting the operators in term of their mean values ($c-$numbers) and some amount of quantum fluctuations as $a_{j}=\alpha_{j}+\delta\alpha_{j}$ and $b_{s_{j}}=\beta_{j}+\delta\beta_{j}$, where $\alpha_{j}=\langle a_{j}\rangle$ and $\beta_{j}=\langle b_{s_{j}}\rangle$. This procedure leads to the classical mean equations,
$\dot{\alpha_{j}}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\alpha_{j}+\sqrt{\kappa}\alpha^{in}_{j},$ (32a)
$\dot{\beta_{j}}$ $=-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\beta_{j}+i\tilde{J}_{m}\beta_{3-j}+i\tilde{g}|\alpha_{j}|^{2},$ (32b)
@@
$\delta\dot{\alpha}_{j}$ $=(i\tilde{\Delta}_{j}-\frac{\kappa}{2})\delta\alpha_{j}+i\tilde{G_{j}}(\delta\beta_{j}+\delta\beta^{\dagger}_{j})+\sqrt{\kappa}\delta\alpha^{in}_{j},$ (33a)
@@
$\rm{\delta\dot{\beta}_{j}}$ $=\rm{-(i\Delta_{m}^{j}+\frac{\gamma_{m}}{2})\delta\beta_{j}}\rm{+i\tilde{J}_{m}\delta\beta_{3-j}}\rm{+i\tilde{G_{j}}(\delta\alpha^{\dagger}_{j}+\delta\alpha_{j})}$
$+\rm{\sqrt{\gamma_{m}}\delta{\beta}^{in}_{j}},$ (33b)
@@
with $\tilde{\Delta}_{j}=\Delta_{j}+2\tilde{g}\Re{\beta_{j}}$, and $\tilde{G_{j}}=\tilde{g}\alpha_{j}$ where the intracavity field amplitude $\alpha_{j}$ is assumed to be real. In order to get an effective mechanical model describing Appendix A, we need to trace out the optical equations. For that purpose, we introduce the following slowly varying operators with tildes, $\delta\alpha_{j}=\delta\tilde{\alpha_{j}}e^{i\tilde{\Delta}_{j}t}$, $\delta\alpha^{in}=\delta\tilde{\alpha^{in}}e^{i\tilde{\Delta}_{j}t}$, and $\delta\beta_{j}=\delta\tilde{\beta_{j}}e^{-i\Delta_{m}^{j}t}$. By using these slowly varying operators in Appendix A, one get
$\dot{\delta\tilde{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}e^{-i(\Delta_{m}^{j}+\tilde{\Delta})t}+\delta\tilde{\beta_{j}}^{\dagger}e^{i(\Delta_{m}^{j}-\tilde{\Delta}_{j})t})$ (34a)
$+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{j},$ (34b)
$\dot{\delta\tilde{\beta_{j}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{i(\Delta_{m}^{j}-\tilde{\Delta}_{j})t}$ (34c)
$+\delta\tilde{\alpha_{j}}e^{i(\Delta_{m}^{j}+\tilde{\Delta}_{j})t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}.$ (34d)
@@
As gain and losses are involved in our system, we assume that the first cavity is driven at the red sideband ($\tilde{\Delta}_{1}=-\Delta_{m}^{j}$) while the second one is driven at the blue sideband ($\tilde{\Delta}_{2}=\Delta_{m}^{j}$). Therefore, on the red sideband we get ($j=1$),
@@
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}+\delta\tilde{\beta_{j}}^{\dagger}e^{2i\Delta_{m}^{j}t})+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{j}},$ (35a)
@@
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{2i\tilde{\Delta}_{j}t}+\delta\tilde{\alpha_{j}})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}},$ (35b)
while on the blue sideband we have ($j=2$),
@@
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}e^{-2i\Delta_{m}^{j}t}+\delta\tilde{\beta_{j}}^{\dagger})+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{j},$ (36a)
@@
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i(\tilde{G_{j}}\delta\tilde{\alpha_{j}}^{\dagger}+\tilde{G_{j}}^{\ast}\delta\tilde{\alpha_{j}}e^{2i\tilde{\Delta}_{j}t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}}.$ (36b)
@@
$\dot{\delta\tilde{\alpha_{1}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{1}}+i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{1},$ (37a)
$\dot{\delta\tilde{\alpha_{2}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{2}}+i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{2},$ (37b)
$\dot{\delta\tilde{\beta_{1}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{1}}+i\tilde{J}_{m}\delta\tilde{\beta}_{2}+i\tilde{G_{1}}^{\ast}\delta\tilde{\alpha_{1}}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{1},$ (37c)
$\dot{\delta\tilde{\beta_{2}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{2}}+i\tilde{J}_{m}\delta\tilde{\beta}_{1}+i\tilde{G_{2}}\delta\tilde{\alpha_{2}}^{\dagger}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{2}.$ (37d)
In order to get the effective mechanical system, we need to trace out the intracavity field ($\delta\tilde{\alpha_{j}}$) in Appendix A. For this purpose, let us introduce $\delta\tilde{\alpha_{j}}=\delta\alpha_{m}^{j}e^{-\frac{\kappa}{2}t}$, which after using them in Appendix A leads to,
$\dot{\delta\alpha_{m}^{1}}$ $=(i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{1}})e^{\frac{\kappa}{2}t},$ (38a)
$\dot{\delta\alpha_{m}^{2}}$ $=(i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{2}})e^{\frac{\kappa}{2}t}.$ (38b)
These equations can be integrated,
@@
where in the weak coupling regime ($\kappa\gg\tilde{G_{j}}$), can be adiabatically integrated and yields,
$\delta\tilde{\alpha}_{1}$ $=\frac{2}{\kappa}(i\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{1}),$ (40a)
$\delta\tilde{\alpha}_{2}$ $=\frac{2}{\kappa}(i\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{in}_{2}).$ (40b)
@@
$\dot{\delta\beta_{1}}$ $=-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})\delta\beta_{1}+i\tilde{J}_{m}\delta\beta_{2}+i\sqrt{\Gamma_{1}^{\ast}}\delta\alpha^{in}+\sqrt{\gamma_{m}}\delta\beta^{in}_{1},$ (41a)
$\dot{\delta\beta_{2}}$ $=-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\delta\beta_{2}+i\tilde{J}_{m}\delta\beta_{1}+i\sqrt{\Gamma_{2}}\delta\alpha^{in\dagger}+\sqrt{\gamma_{m}}\delta\beta^{in}_{2},$ (41b)
where optical induced damping is $\rm{\Gamma}_{j}=\frac{4|\tilde{G_{j}}|^{2}}{\kappa}$ and the effective damping are $\rm{\Gamma_{eff}}^{1,2}=\gamma_{m}\pm\rm{\Gamma}_{1,2}$. From now on, we will assume that $\rm{\Gamma}_{1}=\rm{\Gamma}_{2}=\rm{\Gamma}$. Appendix A can be put in its compact form as,
$\dot{\mathcal{O}}=\mathcal{M}\mathcal{O}+i\sqrt{\mathcal{K}}\mathcal{N}+\sqrt{\gamma_{m}}\mathcal{O}^{in}$ (42)
where $\mathcal{O}\equiv(\delta\beta_{1},\delta\beta_{2})^{T}$, $\mathcal{K}\equiv(\Gamma_{1}^{\ast},\Gamma_{1})^{T}$, $\mathcal{N}\equiv(\delta\alpha^{in},\delta\alpha^{in\dagger})^{T}$ ,$\mathcal{O}^{in}\equiv(\delta\beta^{in}_{1},\delta\beta^{in}_{2})^{T}$, and the matrix
$\mathcal{M}=\begin{bmatrix}-(i\Delta_{m}^{1}+\frac{\rm{\Gamma_{eff}}^{1}}{2})&i\tilde{J}_{m}\\ i\tilde{J}_{m}&-(i\Delta_{m}^{2}+\frac{\rm{\Gamma_{eff}}^{2}}{2})\end{bmatrix}.$ (43)

### 13 Jan 2024 v3

@@ Appendix A Effective mechanical Hamiltonian
$H_{\rm{drive}}$ $={\rm i}E\sum_{j}(a^{\dagger}_{j}-a_{j}),$ (25c)
$H_{\rm{diss}}$ $=-{\rm i}\frac{\kappa}{2}\sum_{j}a^{\dagger}_{j}a_{j}-{\rm i}\frac{\gamma_{m}}{2}\sum_{j}b^{\dagger}_{s_{j}}b_{s_{j}},$ (25d)
where the Hamiltonian $H_{\rm{diss}}$ captures the optical ($\kappa$) and the mechanical ($\gamma_{1}=\gamma_{2}=\gamma_{m}$) dissipations in the system. Furthermore, the driving amplitude has been expressed as $E=\sqrt{\kappa}a^{\rm in}$, where $\alpha^{\rm in}$ is related to the input power $P_{\rm in}$ through $\alpha^{\rm in}=\sqrt{\frac{P_{\rm in}}{\hbar\omega_{p}}}$. The following parameters have been also defined,
$\tilde{g}$ $=g{\rm e}^{-r},$ (26a)
@@
The Quantum Langevin Equations (QLEs) of the Hamiltonian given in Appendix A can be derived using the Heisenberg equation $\dot{\mathcal{O}}={\rm i}[H,\mathcal{O}]+\mathcal{N}$, where $H$ is a Hamiltonian, $\mathcal{O}$ is an operator ($\mathcal{O}\in\{b_{s_{j}},a_{j}\}$, and $\mathcal{N}$ is the quantum noise related to the operator. Following this procedure, the QLEs of our system yield,
$\dot{a}_{j}$ $=-\left({\rm i}[\Delta_{j}+\tilde{g}(b_{s_{j}}+b^{\dagger}_{s_{j}})]+\frac{\kappa}{2}\right)a_{j}+\sqrt{\kappa}\alpha^{\rm in}_{j}+\sqrt{\kappa}a^{\rm in}_{j},$ (27a)
$\dot{b}_{s_{j}}$ $=-\left({\rm i}\Delta_{m}^{j}+\frac{\gamma_{m}}{2}\right)b_{s_{j}}+{\rm i}\tilde{J}_{m}b_{3-j}+{\rm i}\tilde{g}a^{\dagger}_{j}a_{j}+\sqrt{\gamma_{m}}\beta^{\rm in}_{j}.$ (27b)
The nonlinear set of equations displayed in Appendix A can be linearized by splitting the operators in terms of their mean values ($c-$numbers) and some amount of quantum fluctuations as $a_{j}=\alpha_{j}+\delta\alpha_{j}$ and $b_{s_{j}}=\beta_{j}+\delta\beta_{j}$, where $\alpha_{j}\equiv\langle a_{j}\rangle$ and $\beta_{j}\equiv\langle b_{s_{j}}\rangle$. This procedure leads to the classical mean field equations,
$\dot{\alpha_{j}}$ $=-\left({\rm i}\tilde{\Delta}_{j}+\frac{\kappa}{2}\right)\alpha_{j}+\sqrt{\kappa}\alpha^{\rm in}_{j},$ (28a)
$\dot{\beta_{j}}$ $=-\left({\rm i}\Delta_{m}^{j}+\frac{\gamma_{m}}{2}\right)\beta_{j}+{\rm i}\tilde{J}_{m}\beta_{3-j}+{\rm i}\tilde{g}|\alpha_{j}|^{2}$ (28b)
@@
$\delta\dot{\alpha}_{j}$ $=-\left({\rm i}\tilde{\Delta}_{j}+\frac{\kappa}{2}\right)\delta\alpha_{j}+{\rm i}\tilde{G_{j}}(\delta\beta_{j}+\delta\beta^{\dagger}_{j})+\sqrt{\kappa}\delta\alpha^{\rm in}_{j},$ (29a)
@@
$\rm{\delta\dot{\beta}_{j}}$ $=-\left({\rm i}\Delta_{m}^{j}+\frac{\gamma_{m}}{2}\right)\delta\beta_{j}+{\rm i}\tilde{J}_{m}\delta\beta_{3-j}+{\rm i}\tilde{G_{j}}(\delta\alpha^{\dagger}_{j}+\delta\alpha_{j})$
$+\sqrt{\gamma_{m}}\delta{\beta}^{\rm in}_{j},$ (29b)
@@
with $\tilde{\Delta}_{j}=\Delta_{j}+2\tilde{g}{\rm Re}(\beta_{j})$, and $\tilde{G_{j}}=\tilde{g}\alpha_{j}$ where the intracavity field amplitude $\alpha_{j}$ is assumed to be real. In order to get an effective mechanical model describing Appendix A, we need to trace out the optical equations. For that purpose, we introduce the following slowly varying operators with tildes, $\delta\alpha_{j}=\delta\tilde{\alpha_{j}}{\rm e}^{{\rm i}\tilde{\Delta}_{j}t}$, $\delta\alpha^{\rm in}=\delta\tilde{\alpha^{\rm in}}{\rm e}^{{\rm i}\tilde{\Delta}_{j}t}$, and $\delta\beta_{j}=\delta\tilde{\beta_{j}}{\rm e}^{-{\rm i}\Delta_{m}^{j}t}$. By using these slowly varying operators in Appendix A, one gets
$\dot{\delta\tilde{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+{\rm i}\tilde{G_{j}}(\delta\tilde{\beta_{j}}{\rm e}^{-{\rm i}(\Delta_{m}^{j}+\tilde{\Delta})t}+\delta\tilde{\beta_{j}}^{\dagger}{\rm e}^{{\rm i}(\Delta_{m}^{j}-\tilde{\Delta}_{j})t})$ (30a)
$+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{j},$ (30b)
$\dot{\delta\tilde{\beta_{j}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+{\rm i}\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+{\rm i}\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}{\rm e}^{{\rm i}(\Delta_{m}^{j}-\tilde{\Delta}_{j})t}$ (30c)
$+\delta\tilde{\alpha_{j}}{\rm e}^{{\rm i}(\Delta_{m}^{j}+\tilde{\Delta}_{j})t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{\rm in}_{j}.$ (30d)
@@
As gain and losses are involved in our system, we assume that the first cavity is driven at the red sideband ($\tilde{\Delta}_{1}=-\Delta_{m}^{j}$) while the second one is driven at the blue sideband ($\tilde{\Delta}_{2}=\Delta_{m}^{j}$). Therefore, on the red sideband we get ($j=1$):
@@
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+{\rm i}\tilde{G_{j}}(\delta\tilde{\beta_{j}}+\delta\tilde{\beta_{j}}^{\dagger}{\rm e}^{2i\Delta_{m}^{j}t})+\sqrt{\kappa}\delta\tilde{\alpha^{\rm in}_{j}},$ (31a)
@@
$\delta\tilde{\dot{\beta_{j}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+{\rm i}\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+{\rm i}\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}{\rm e}^{2{\rm i}\tilde{\Delta}_{j}t}+\delta\tilde{\alpha_{j}})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{\rm in}_{j},$ (31b)
while on the blue sideband we have ($j=2$):
@@
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+{\rm i}\tilde{G_{j}}(\delta\tilde{\beta_{j}}{\rm e}^{-2i\Delta_{m}^{j}t}+\delta\tilde{\beta_{j}}^{\dagger})+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{j},$ (32a)
@@
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+{\rm i}\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+{\rm i}(\tilde{G_{j}}\delta\tilde{\alpha_{j}}^{\dagger}+\tilde{G_{j}}^{\ast}\delta\tilde{\alpha_{j}}{\rm e}^{2i\tilde{\Delta}_{j}t})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{\rm in}_{j}}.$ (32b)
@@
$\dot{\delta\tilde{\alpha_{1}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{1}}+{\rm i}\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{1},$ (33a)
$\dot{\delta\tilde{\alpha_{2}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{2}}+{\rm i}\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{2},$ (33b)
$\dot{\delta\tilde{\beta_{1}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{1}}+{\rm i}\tilde{J}_{m}\delta\tilde{\beta}_{2}+{\rm i}\tilde{G_{1}}^{\ast}\delta\tilde{\alpha_{1}}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{\rm in}_{1},$ (33c)
$\dot{\delta\tilde{\beta_{2}}}$ $=-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{2}}+{\rm i}\tilde{J}_{m}\delta\tilde{\beta}_{1}+{\rm i}\tilde{G_{2}}\delta\tilde{\alpha_{2}}^{\dagger}+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{\rm in}_{2}.$ (33d)
In order to get the effective mechanical system, we need to trace out the intracavity field ($\delta\tilde{\alpha_{j}}$) in Appendix A. For this purpose, let us introduce $\delta\tilde{\alpha_{j}}=\delta\alpha_{m}^{j}{\rm e}^{-\frac{\kappa}{2}t}$, which after using them in Appendix A leads to,
$\dot{\delta\alpha_{m}^{1}}$ $=({\rm i}\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha^{\rm in}_{1}}){\rm e}^{\frac{\kappa}{2}t},$ (34a)
$\dot{\delta\alpha_{m}^{2}}$ $=({\rm i}\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha^{\rm in}_{2}}){\rm e}^{\frac{\kappa}{2}t}.$ (34b)
These equations can be integrated as follows
@@
where in the weak coupling regime ($\kappa\gg\tilde{G_{j}}$), can be adiabatically integrated and yields:
$\delta\tilde{\alpha}_{1}$ $=\frac{2}{\kappa}({\rm i}\tilde{G_{1}}\delta\tilde{\beta_{1}}+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{1}),$ (36a)
$\delta\tilde{\alpha}_{2}$ $=\frac{2}{\kappa}({\rm i}\tilde{G_{2}}\delta\tilde{\beta_{2}}^{\dagger}+\sqrt{\kappa}\delta\tilde{\alpha}^{\rm in}_{2}).$ (36b)
@@
$\dot{\delta\beta_{1}}$ $=-({\rm i}\Delta_{m}^{1}+\frac{\rm{\Gamma_{\rm eff}}^{1}}{2})\delta\beta_{1}+{\rm i}\tilde{J}_{m}\delta\beta_{2}+{\rm i}\sqrt{\Gamma_{1}^{\ast}}\delta\alpha^{\rm in}+\sqrt{\gamma_{m}}\delta\beta^{\rm in}_{1},$ (37a)
$\dot{\delta\beta_{2}}$ $=-({\rm i}\Delta_{m}^{2}+\frac{\rm{\Gamma_{\rm eff}}^{2}}{2})\delta\beta_{2}+{\rm i}\tilde{J}_{m}\delta\beta_{1}+{\rm i}\sqrt{\Gamma_{2}}\delta\alpha^{in\dagger}+\sqrt{\gamma_{m}}\delta\beta^{\rm in}_{2},$ (37b)
where optical induced damping is $\rm{\Gamma}_{j}:=\frac{4|\tilde{G_{j}}|^{2}}{\kappa}$ and the effective damping are $\rm{\Gamma^{1,2}_{\rm eff}}=\gamma_{m}\pm\rm{\Gamma}_{1,2}$. From now on, we will assume that $\rm{\Gamma}_{1}=\rm{\Gamma}_{2}\equiv\rm{\Gamma}$. Appendix A can be put in its compact form as
$\dot{\mathcal{O}}=\mathcal{M}\mathcal{O}+{\rm i}\sqrt{\mathcal{K}}\mathcal{N}+\sqrt{\gamma_{m}}\mathcal{O}^{\rm in}$ (38)
where $\mathcal{O}:=(\delta\beta_{1},\delta\beta_{2})^{\top}$, $\mathcal{K}:=(\Gamma_{1}^{\ast},\Gamma_{1})^{\top}$, $\mathcal{N}:=(\delta\alpha^{\rm in},\delta\alpha^{in\dagger})^{\top}$ ,$\mathcal{O}^{\rm in}:=(\delta\beta^{\rm in}_{1},\delta\beta^{\rm in}_{2})^{\top}$, and the matrix
$\mathcal{M}:=\begin{bmatrix}-\left({\rm i}\Delta^{1}_{m}+\frac{\rm{\Gamma^{1}_{\rm eff}}}{2}\right)&{\rm i}\tilde{J}_{m}\\ {\rm i}\tilde{J}_{m}&-\left({\rm i}\Delta_{m}^{2}+\frac{\rm{\Gamma^{2}_{\rm eff}}}{2}\right)\end{bmatrix}.$ (39)